A boat travels 60 km upstream, (against the current) in 5 hours. The boat travels the same distance downstream in 3 hours. What is the rate of the boat in still water?


Answer by checkley77(12569)

60/5=12 kph.
60/3=20 kph.
(20-12)/2=8/2=4 kph is the speed of the current.
12+4=16 kph is the speed of the boat in still water.
20-4=16 kph. ditto.